Homology Stability for Unitary Groups

B. Mirzaii, W. van der Kallen

Received: January 29, 2002

Communicated by Ulf Rehmann

Abstract. In this paper the homology stability for unitary groups over a ring with finite unitary stable rank is established. First we develop a ‘nerve theorem’ on the homotopy type of a poset in terms of a cover by subposets, where the cover is itself indexed by a poset. We use the nerve theorem to show that a poset of sequences of isotropic vectors is highly connected, as conjectured by Charney in the eighties. Homology stability of symplectic groups and orthogonal groups appear as a special case of our results.

2000 Mathematics Subject Classification: Primary 19G99; Secondary 11E70, 18G30, 19B10. Keywords and Phrases: Poset, acyclicity, unitary groups, homology stability.

1. Introduction

Interest in homological stability problems in algebraic K-theory started with Quillen, who used it in [15] to study the higher K-groups of a ring of integers. As a result of stability he proved that these groups are finitely generated (see also [7]). After that there has been considerable interest in homological stability for general linear groups. The most general results in this direction are due to the second author [20] and Suslin [19]. Parallel to this, similar questions for other classical groups such as orthogonal and symplectic groups have been studied. For work in this direction, see [23], [1], [5], [12], [13]. The most general result is due to Charney [5]. She proved the homology stability for orthogonal and symplectic groups over a Dedekind domain. Panin in [13] proved a similar result but with a different method and with better range of stability. Our goal in this paper is to prove that homology stabilizes of the unitary groups over rings with finite unitary stable rank. To do so we prove that the poset of isotropic unimodular sequences is highly connected. Recall that Panin in [12] had already sketched how one can do this for a finite dimensional affine algebra over an infinite field, in the case of symplectic and orthogonal groups. However, while the assumption about the infinite field provides a significant simplification, it excludes cases of primary interest, namely rings that are finitely generated over the integers. Our approach is as follows. We first extend a theorem of Quillen [16, Thm 9.1] which was his main tool to prove that certain posets are highly connected. We use it to develop a quantitative analogue for posets of the nerve theorem, which expresses the homotopy type of a space in terms of the the nerve of a suitable cover. In our situation both the elements of the cover and the nerve are replaced with posets. We work with posets of ordered sequences ‘satisfying the chain condition’, as this is a good replacement for simplicial complexes in the presence of group actions. (Alternatively one might try to work with barycentric subdivisions of a simplicial complex.) The new nerve theorem allows us to exploit the higher connectivity of the poset of unimodular sequences due to the second author. The higher connectivity of the poset of isotropic unimodular sequences follows inductively. We conclude with the homology stability theorem.

2. Preliminaries

Recall that a topological space X is (−1)-connected if it is non-empty, 0connected if it is non-empty and path connected, 1-connected if it is non-empty and simply connected. In general for n ≥ 1, X is called n-connected if X is nonempty, X is 0-connected and π_{i}(X, x) = 0 for every base point x ∈ X and 1 ≤ i ≤ n. For n ≥ −1 a space X is called n-acyclic if it is nonempty and H̃_{i}(X, ℤ) = 0 for 0 ≤ i ≤ n. For n < −1 the conditions of n-connectedness and n-acyclicness are vacuous.

Theorem 2.1 (Hurewicz). For n ≥ 0, a topological space X is n-connected if and only if the reduced homology groups H̃_{i}(X, ℤ) are trivial for 0 ≤ i ≤ n and X is 1-connected if n ≥ 1.

Proof. See [25], Chap. IV, Corollaries 7.7 and 7.8. □

Let X be a partially ordered set or briefly a poset. Consider the simplicial complex associated to X, that is the simplicial complex where vertices or 0simplices are the elements of X and the k-simplices are the (k + 1)-tuples (x_{0}, …, x_{k}) of elements of X with x_{0} < ⋯ < x_{k}. We denote it again by X. We denote the geometric realization of X by |X| and we consider it with the weak topology. It is well known that |X| is a CW-complex [11]. By a morphism or map of posets f : X → Y we mean an order-preserving map i. e. if x ≤ x′ then f(x) ≤ f(x′). Such a map induces a continuous map |f| : |X| → |Y |.

Remark 2.2. If K is a simplicial complex and X the partially ordered set of simplices of K, then the space |X| is the barycentric subdivision of K. Thus every simplicial complex, with weak topology, is homeomorphic to the geometric realization of some, and in fact many, posets. Furthermore since it is well known that any CW-complex is homotopy equivalent to a simplicial complex, it follows that any interesting homotopy type is realized as the geometric realization of a poset.

Proposition 2.3. Let X and Y be posets.

(i) (Segal [17]) If f, g : X → Y are maps of posets such that f(x) ≤ g(x) for all x ∈ X, then |f| and |g| are homotopic. (ii) If the poset X has a minimal or maximal element then |X| is contractible. (iii) If X^{op} denotes the opposite poset of X, i. e. with opposite ordering, then |X^{op}| ≃ |X|.

Proof. (i) Consider the poset I = {0, 1 : 0 < 1} and define the poset map h : I × X → Y as h(0, x) = f(x), h(1, x) = g(x). Since |I| ≃ [0, 1], we have |h| : [0, 1] × |X| → |Y | with |h|(0, x) = |f|(x) and |h|(1, x) = |g|(x). This shows that |f| and |g| are homotopic. (ii) Suppose X has a maximal element z. Consider the map f : X → X with f(x) = z for every x ∈ X. Clearly for every x ∈ X, id_{X}(x) ≤ f(x). This shows that id_{X} and the constant map f are homotopic. So X is contractible. If X has a minimal element the proof is similar. (iii). This is natural and easy. □

The construction X ↦ |X| allows us to assign topological concepts to posets. For example we define the homology groups of a poset X to be those of |X|, we call X n-connected or contractible if |X| is n-connected or contractible etc. Note that X is connected if and only if X is connected as a poset. By the dimension of a poset X, we mean the dimension of the space |X|, or equivalently the supremum of the integers n such that there is a chain x_{0} < ⋯ < x_{n} in X. By convention the empty set has dimension −1. Let X be a poset and x ∈ X. Define Link_{X}^{+}(x) := {u ∈ X : u > x} and Link_{X}^{−}(x) := {u ∈ X : u < x}. Given a map f : X → Y of posets and an element y ∈ Y, define subposets f/y and y\f of X as follows

f/y := {x ∈ X : f(x) ≤ y} y\f := {x ∈ X : f(x) ≥ y}.

In fact f/y = f^{−1}(Y_{≤y}) and y\f = f^{−1}(Y_{≥y}) where Y_{≤y} = {z ∈ Y : z ≤ y} and Y_{≥y} = {z ∈ Y : z ≥ y}. Note that by 2.3 (ii), Y_{≤y} and Y_{≥y} are contractible. If id_{Y} : Y → Y is the identity map, then id_{Y} /y = Y_{≤y} and y\id_{Y} = Y_{≥y}. Let F : X → \underline{Ab} be a functor from a poset X, regarded as a category in the usual way, to the category of abelian groups. We define the homology groups H_{i}(X, F) of X with coefficient F to be the homology of the complex C_{∗}(X, F) given by

C_{n}(X, F) = ⨁_{x_{0}<···<x_{n}}F(x_{0})

where the direct sum is taken over all n-simplices in X, with differential ∂_{n} = Σ_{i=0}^{n}(−1)^{i}d_{i}^{n} where d_{i}^{n} : C_{n}(X, F) → C_{n−1}(X, F) and d_{i}^{n} takes the (x_{0} < ⋯ < x_{n})-component of C_{n}(X, F) to the (x_{0} < ⋯ < x̂_{i} < ⋯ < x_{n})-component of C_{n−1}(X, F) via d_{i}^{n} = id_{F(x_{0})} if i > 0 and d_{0}^{n} : F(x_{0}) → F(x_{1}). In particular, for the empty set we have H_{i}(∅, F) = 0 for i ≥ 0.

Let F be the constant functor ℤ. Then the homology groups with this coefficient coincide with the integral homology of |X|, that is H_{k}(X, ℤ) = H_{k}(|X|, ℤ) for all k ∈ ℤ, [6, App. II]. Let H̃_{i}(X, ℤ) denote the reduced integral homology of the poset X, that is H̃_{i}(X, ℤ) = ker{H_{i}(X, ℤ) → H_{i}(pt, ℤ)} if X ≠ ∅ and

H̃_{i}(∅, ℤ) = \begin{cases}ℤ & if i = −1 \\ 0 & if i ≠ −1 \end{cases}. So H̃_{i}(X, ℤ) = H_{i}(X, ℤ) for i ≥ 1 and for i = 0 we

have the exact sequence

0 → H̃_{0}(X, ℤ) → H_{0}(X, ℤ) → ℤ → H̃_{−1}(X, ℤ) → 0

where ℤ is identified with the group H_{0}(pt, ℤ). Notice that H_{0}(X, ℤ) is identified with the free abelian group generated by the connected components of X. A local system of abelian groups on a space (resp. poset) X is a functor F from the groupoid of X (resp. X viewed as a category), to the category of abelian groups which is morphism-inverting, i. e. such that the map F(x) → F(x′) associated to a path from x to x′ (resp. x ≤ x′) is an isomorphism. Clearly, a local system F on a path connected space (resp. 0-connected poset) is determined, up to canonical isomorphism, by the following data: if x ∈ X is a base point, it suffices to be given the group F(x) and an action of π_{1}(X, x) on F(x). The homology groups H_{k}(X, F) of a space X with a local system F are a generalization of the ordinary homology groups. In fact if X is a 0-connected space and if F is a constant local system on X, then H_{k}(X, F) ≃ H_{k}(X, F(x_{0})) for every x_{0} ∈ X [25, Chap. VI, 2.1]. Let X be a poset and F a local system on |X|. Then the restriction of F to X is a local system on X. Considering F as a functor from X to the category of abelian groups, we can define H_{k}(X, F) as in the above. Conversely if F is a local system on the poset X, then there is a unique local system, up to isomorphism, on |X| such that the restriction to X is F [25, Chap. VI, Thm 1.12], [14, I, Prop. 1]. We denote both local systems by F.

Theorem 2.4. Let X be a poset and F a local system on X. Then the homology groups H_{k}(|X|, F) are isomorphic with the homology groups H_{k}(X, F).

Proof. See [25, Chap. VI, Thm. 4.8] or [14, I, p. 91]. □

Theorem 2.5. Let X be a path connected space with a base point x and let F be a local system on X. Then the inclusion {x} ↪ X induces an isomorphism F(x)/G ⟶^{≃} H_{0}(X, F) where G is the subgroup of F(x) generated by all the elements of the form a − βa with a ∈ F(x), β ∈ π_{1}(X, x).

Proof. See [25], Chap. VI, Thm. 2.8^{∗} and Thm. 3.2. □

We need the following interesting and well known lemma about the covering spaces of the space |X|, where X is a poset (or more generally a simplicial set). For a definition of a covering space, useful for our purpose, and some more information, see [18, Chap. 2].

Lemma 2.6. For a poset X the category of the covering spaces of the space |X| is equivalent to the category L_{S}(X), the category of functors F : X → \underline{Set}, where \underline{Set} is the category of sets, such that F(x) → F(x′) is a bijection for every relation x ≤ x′.

Proof. See [16, Section 7] or [14, I, p. 90]. □

3. Homology and homotopy of posets

Theorem 3.1. Let f : X → Y be a map of posets. Then there is a first quadrant spectral sequence

E_{p,q}^{2} = H_{p}(Y, y ↦ H_{q}(f/y, ℤ)) ⇒ H_{p+q}(X, ℤ).

The spectral sequence is functorial, in the sense that if there is a commutative diagram of posets

X′ ⟶^{f′} Y′
↓^{g_{X}}  ↓^{g_{Y}}
X ⟶^{f} Y

then there is a natural map from the spectral sequence arising from f′ to the spectral sequence arising from f. Moreover the map g_{X}_{∗} : H_{i}(X′, ℤ) → H_{i}(X, ℤ) is compatible with this natural map.

Proof. Let C_{∗,∗}(f) be the double complex such that C_{p,q}(f) is the free abelian group generated by the set {(x_{0} < ⋯ < x_{q}, f(x_{q}) < y_{0} < ⋯ < y_{p}) : x_{i} ∈ X, y_{i} ∈ Y }. The first spectral sequence of this double complex has as E^{1}- term E_{p,q}^{1}(I) = H_{q}(C_{p,∗}(f)) = ⨁_{y_{0}<···<y_{p}} H_{q}(f/y_{0}, ℤ). By the general theory of double complexes (see for example [24, Chap. 5]), we know that E_{p,q}^{2}(I) is the homology of the chain complex E_{∗,q}^{1}(I) = C_{∗}(Y, G_{q}) where G_{q} : Y → \underline{Ab}, G_{q}(y) = H_{q}(f/y, ℤ) and hence E_{p,q}^{2}(I) = H_{p}(Y, G_{q}) = H_{p}(Y, y ↦ H_{q}(f/y, ℤ)). The second spectral sequence has as E^{1}-term E_{p,q}^{1}(II) = H_{q}(C_{∗,p}(f)) = ⨁_{f(x_{p})<y_{0}<···<y_{q}} H_{q}(f(x_{p})\id_{Y} , ℤ). But by 2.3 (ii), f(x_{p})\id_{Y} = Y_{≥f(x_{p})} is contractible, so E_{∗,0}^{1}(II) = C_{∗}(X^{op}, ℤ) and E_{∗,q}^{1}(II) = 0 for q > 0. Hence H_{i}(Tot(C_{∗,∗}(f))) ≃ H_{i}(X^{op}, ℤ) ≃ H_{i}(X, ℤ). This completes the proof of existence and convergence of the spectral sequence. The functorial behavior of the spectral sequence follows from the functorial behavior of the spectral sequence of a filtration [24, 5.5.1] and the fact that the first and the second spectral sequences of the double complex arise from some filtrations. □

Remark 3.2. The above spectral sequence is a special case of a more general Theorem [6, App. II]. The above proof is taken from [9, Chap. I] where the functorial behavior of the spectral sequence is more visible. For more details see [9].

Definition 3.3. Let X be a poset. A map ht_{X} : X → ℤ_{≥0} is called height function if it is a strictly increasing map.

Example 3.4. The height function ht_{X}(x) = 1 + dim(Link_{X}^{−}(x)) is the usual one considered in [16], [9] and [5].

Lemma 3.5. Let X be a poset such that Link_{X}^{+}(x) is (n − ht_{X}(x) − 2)-acyclic, for every x ∈ X, where ht_{X} is a height function on X. Let F : X → \underline{Ab} be a functor such that F(x) = 0 for all x ∈ X with ht_{X}(x) ≥ m, where m ≥ 1. Then H_{k}(X, F) = 0 for k ≤ n − m.

Proof. First consider the case of a functor F such that F(x) = 0 if ht_{X}(x) ≠

m − 1. Then C_{k}(X, F) = ⨁_{x_{0}<···<x_{k}_{ht_{X} (x_{0})=m−1}}F(x_{0}). Clearly 0 = d_{0}^{k} = F(x_{0} < x_{1}) =

F(x_{0}) → F(x_{1}). Thus ∂_{k} = Σ_{i=1}^{k}(−1)^{i}d_{i}^{k}. Define C_{−1}(Link_{X}^{+}(x_{0}), F(x_{0})) = F(x_{0}) and complete the singular complex of Link_{X}^{+}(x_{0}) with coefficient in F(x_{0}) to

⋯ → C_{0}(Link_{X}^{+}(x_{0}), F(x_{0})) →^{ε} C_{−1}(Link_{X}^{+}(x_{0}), F(x_{0})) → 0

where ε((g_{i})) = Σ_{i}g_{i}. Then

C_{k}(X, F) = ⨁_{ht_{X}(x_{0})=m−1}( ⨁_{x_{1}<···<x_{k}_{x_{0}<x_{1}}}F(x_{0}))

= ⨁_{ht_{X}(x_{0})=m−1}C_{k−1}(Link_{X}^{+}(x_{0}), F(x_{0})).

The complex C_{k−1}(Link_{X}^{+}(x_{0}), F(x_{0})) is the standard complex for computing the reduced homology of Link_{X}^{+}(x_{0}) with constant coefficient F(x_{0}). So

H_{k}(X, F) = ⨁_{ht_{X}(x)=m−1}H̃_{k−1}(Link_{X}^{+}(x), F(x)).

If ht_{X}(x_{0}) = m−1 then Link_{X}^{+}(x_{0}) is (n−(m−1)−2)-acyclic, and by the universal coefficient theorem [18, Chap. 5, Thm. 8], H̃_{k−1}(Link_{X}^{+}(x_{0}), F(x_{0})) = 0 for −1 ≤ k−1 ≤ n−(m−1)−2. This shows that H_{k}(X, F) = 0 for 0 ≤ k ≤ n−m. To prove the lemma in general, we argue by induction on m. If m = 1 then for ht_{X}(x) ≥ 1, F(x) = 0. So the lemma follows from the special case above. Suppose m ≥ 2. Define F_{0} and F_{1} to be the functors

F_{0}(x) = \begin{cases}F(x) & if ht_{X}(x) < m − 1 \\ 0 & if ht_{X}(x) ≥ m − 1 \end{cases} , F_{1}(x) = \begin{cases}F(x) & if ht_{X}(x) = m − 1 \\ 0 & if ht_{X}(x) ≠ m − 1 \end{cases}

respectively. Then there is a short exact sequence 0 → F_{1} → F → F_{0} → 0. By the above discussion, H_{k}(X, F_{1}) = 0 for 0 ≤ k ≤ n − m and by induction for m − 1, we have H_{k}(X, F_{0}) = 0 for k ≤ n − (m − 1). By the long exact sequence for the above short exact sequence of functors it is easy to see that H_{k}(X, F) = 0 for 0 ≤ k ≤ n − m. □

Theorem 3.6. Let f : X → Y be a map of posets and ht_{Y} a height function on Y. Assume for every y ∈ Y, that Link_{Y}^{+}(y) is (n − ht_{Y} (y) − 2)-acyclic and f/y is (ht_{Y} (y) − 1)-acyclic. Then f_{∗} : H_{k}(X, ℤ) → H_{k}(Y, ℤ) is an isomorphism for 0 ≤ k ≤ n − 1.

Proof. By theorem 3.1, we have the first quadrant spectral sequence

E_{p,q}^{2} = H_{p}(Y, y ↦ H_{q}(f/y, ℤ)) ⇒ H_{p+q}(X, ℤ).

Since H_{q}(f/y, ℤ) = 0 for 0 < q ≤ ht_{Y} (y) − 1, the functor G_{q} : Y → \underline{Ab}, G_{q}(y) = H_{q}(f/y, ℤ) is trivial for ht_{Y} (y) ≥ q + 1, q > 0. By lemma 3.5, H_{p}(Y, G_{q}) = 0 for p ≤ n − (q + 1). Hence E_{p,q}^{2} = 0 for p + q ≤ n − 1, q > 0. If q = 0, by writing the long exact sequence for the short exact sequence 0 → H̃_{0}(f/y, ℤ) → H_{0}(f/y, ℤ) → ℤ → 0, valid because f/y is nonempty, we have

⋯ → H_{n}(Y, ℤ) → H_{n−1}(Y, y ↦ H̃_{0}(f/y, ℤ)) → E_{n−1,0}^{2} →
⋯ → H_{1}(Y, ℤ) → H_{0}(Y, y ↦ H̃_{0}(f/y, ℤ)) → E_{0,0}^{2} → H_{0}(Y, ℤ) → 0.

If ht_{Y} (y) ≥ 1, then H̃_{0}(f/y, ℤ) = 0. By lemma 3.5, H_{k}(Y, y ↦ H̃_{0}(f/y, ℤ)) = 0 for 0 ≤ k ≤ n − 1. Thus

E_{p,q}^{2} = \begin{cases}H_{p}(Y, ℤ) & if q = 0, 0 ≤ p ≤ n − 1 \\ 0 & if p + q ≤ n − 1, q > 0 \end{cases} .

This shows that E_{p,q}^{2} ≃ ⋯ ≃ E_{p,q}^{∞} for 0 ≤ p + q ≤ n − 1. Therefore H_{k}(X, ℤ) ≃ H_{k}(Y, ℤ) for 0 ≤ k ≤ n − 1. Now consider the commutative diagram

\begin{matrix}X & ⟶^{f} & Y \\ ↓^{f} & & ↓^{id}_{Y} \\ Y & ⟶^{id_{Y}} & Y \end{matrix} .

By functoriality of the spectral sequence 3.1, and the above calculation we get the diagram

\begin{aligned}H_{k}(Y, y ↦ H_{0}(f/y, ℤ)) & ⟶^{≃} & H_{k}(X, ℤ) \\ ↓^{id}_{Y} _{∗} & & ↓^{f_{∗}} \\ H_{k}(Y, y ↦ H_{0}(id_{Y} /y, ℤ)) & ⟶^{≃} & H_{k}(Y, ℤ) \end{aligned} .

Since id_{Y} /y = Y_{≤y} is contractible, we have H_{k}(Y, y ↦ H_{0}(id_{Y} /y, ℤ)) = H_{k}(Y, ℤ). The map id_{Y} _{∗} is an isomorphism for 0 ≤ k ≤ n − 1, from the above long exact sequence. This shows that f_{∗} is an isomorphism for 0 ≤ k ≤ n − 1. □

Lemma 3.7. Let X be a 0-connected poset. Then X is 1-connected if and only if for every local system F on X and every x ∈ X, the map F(x) → H_{0}(X, F), induced from the inclusion {x} ↪ X, is an isomorphism (or equivalently, every local system on X is a isomorphic with a constant local system).

Proof. If X is 1-connected then by theorem 2.5 and the connectedness of X, one has F(x) ⟶^{≃} H_{0}(X, F) for every x ∈ X. Now let every local system on X be isomorphic with a constant local system. Let F : X → \underline{Set} be in L_{S}(X).

Define the functor G : X → \underline{Ab} where G(x) is the free abelian group generated by F(x). Clearly G is a local system and so it is constant system. It follows that F is isomorphic to a constant functor. So by lemma 2.6, any connected covering space of |X| is isomorphic to |X|. This shows that the universal covering of |X|, is |X|. Note that the universal covering of a connected simplicial simplex exists and is simply connected [18, Chap. 2, Cor. 14 and 15]. Therefore X is 1-connected. □

Theorem 3.8. Let f : X → Y be a map of posets and ht_{Y} a height function on Y. Assume for every y ∈ Y, that Link_{Y}^{+}(y) is (n − ht_{Y} (y) − 2)-connected and f/y is (ht_{Y} (y) − 1)-connected. Then X is (n − 1)-connected if and only if Y is (n − 1)-connected.

Proof. By 2.1 and 3.6 we may assume n ≥ 2. So it is enough to prove that X is 1-connected if and only if Y is 1-connected. Let F : X → \underline{Ab} be a local system. Define the functor G : Y → \underline{Ab} with

G(y) = \begin{cases}H_{0}(f/y, F) & if ht_{Y} (y) ≠ 0 \\ H_{0}(Link_{Y}^{+}(y), y′ ↦ H_{0}(f/y′, F)) & if ht_{Y} (y) = 0 \end{cases} .

We prove that G is a local system. If ht_{Y} (y) ≥ 2 then f/y is 1-connected and by 3.6, F|_{f/y} is a constant system, so by 3.7, H_{0}(f/y, F) ≃ F(x) for every x ∈ f/y. If ht_{Y} (y) = 1, then f/y is 0-connected and Link_{Y}^{+}(y) is nonempty. Choose y′ ∈ Y such that y < y′. Now f/y′ is 1-connected and so F|_{f/y′} is a constant system on f/y′. But f/y ⊂ f/y′, so F|_{f/y} is a constant system. Since f/y is 0-connected, by 2.5 and the fact that we mentioned before theorem 2.4, H_{0}(f/y, F) ≃ F(x) for every x ∈ f/y. Now let ht_{Y} (y) = 0. Then Link_{Y}^{+}(y) is 0-connected, f/y is nonempty and for every y′ ∈ Link_{Y}^{+}(y), H_{0}(f/y′, F) ≃ H_{0}((f/y)^{∘}, F) where (f/y)^{∘} is a component of f/y, which we fix. This shows that the local system F′ : Link_{Y}^{+}(y) → \underline{Ab} with y′ ↦ H_{0}(f/y′, F) is isomorphic to a constant system, so H_{0}(Link_{Y}^{+}(y), y′ ↦ H_{0}(f/y′, F)) = H_{0}(Link_{Y}^{+}(y), F′) ≃ F′(y′) ≃ F(x) for every x ∈ f/y′. Therefore G is a local system. If Y is 1-connected, by 3.7, G is a constant system. But it is easy to see that F ≃ G ∘ f. Therefore F is a constant system. Since X is connected by our homology calculation, by 3.7 we conclude that X is 1-connected. Now let X be 1-connected. If E is a local system on Y, then f^{∗}E := E ∘ f is a local system on X. So it is a constant local system. As above we can construct a local system G′ on Y from F′ := E ∘ f. This gives a natural transformation from G′ to E which is an isomorphism. Since E ∘ f is constant, by 2.5 and 3.7 and an argument as above one sees that G′ is constant. Therefore E is isomorphic to a constant local system and 3.7 shows that Y is 1-connected. □

Remark 3.9. In the proof of the above theorem 3.8 we showed in fact that: Let f : X → Y be a map of posets and ht_{Y} a height function on Y and n ≥ 2. Assume for every y ∈ Y, that Link_{Y}^{+}(y) is (n − ht_{Y} (y) − 2)-connected and f/y is (ht_{Y} (y) − 1)-connected. Then f^{∗} : L_{S}(Y) → L_{S}(X), with E ↦ E ∘ f is an equivalence of categories.

Remark 3.10. Theorem 3.8 is a generalization of a theorem of Quillen [16, Thm. 9.1]. We proved that the converse of that theorem is also valid. Our proof is similar in outline to the proof by Quillen. Furthermore, lemma 3.5 is a generalized version of lemma 1.3 from [5]. With more restrictions, Maazen, in [9, Chap. II] gave an easier proof of Quillen’s theorem.

4. Homology and homotopy of posets of sequences

Let V be a nonempty set. We denote by O(V) the poset of finite ordered sequences of distinct elements of V, the length of each sequence being at least one. The partial ordering on O(V) is defined by refinement: (v_{1}, …, v_{m}) ≤ (w_{1}, …, w_{n}) if and only if there is a strictly increasing map ϕ : {1, …, m} → {1, …, n} such that v_{i} = w_{ϕ(i)}, in other words, if (v_{1}, …, v_{m}) is an order preserving subsequence of (w_{1}, …, w_{n}). If v = (v_{1}, …, v_{m}) we denote by |v| the length of v, that is |v| = m. If v = (v_{1}, …, v_{m}) and w = (w_{1}, …, w_{n}), we write (v_{1}, …, v_{m}, w_{1}, …, w_{n}) as vw. For v ∈ F, but for such v only, we define F_{v} to be the set of w ∈ F such that wv ∈ F. Note that (F_{v})_{w} = F_{wv}. A subset F of O(V) is said to satisfy the chain condition if v ∈ F whenever w ∈ F, v ∈ O(V) and v ≤ w. The subposets of O(V) which satisfy the chain condition are extensively studied in [9], [20] and [4]. In this section we will study them some more. Let F ⊆ O(V). For a nonempty set S we define the poset F ⟨S⟩ as

F ⟨S⟩ := {((v_{1}, s_{1}), …, (v_{r}, s_{r})) ∈ O(V × S) : (v_{1}, …, v_{r}) ∈ F }.

Assume s_{0} ∈ S and consider the injective poset map l_{s_{0}} : F → F ⟨S⟩ with (v_{1}, …, v_{r}) ↦ ((v_{1}, s_{0}), …, (v_{r}, s_{0})). We have clearly a projection p : F ⟨S⟩ → F with ((v_{1}, s_{1}), …, (v_{r}, s_{r})) ↦ (v_{1}, …, v_{r}) such that p ∘ l_{s_{0}} = id_{F}.

Lemma 4.1. Suppose F ⊆ O(V) satisfies the chain condition and S is a nonempty set. Assume for every v ∈ F, that F_{v} is (n − |v|)-connected.

(i) If s_{0} ∈ S then (l_{s_{0}})_{∗} : H_{k}(F, ℤ) → H_{k}(F ⟨S⟩, ℤ) is an isomorphism for 0 ≤ k ≤ n. (ii) If F is min{1, n − 1}-connected, then (l_{s_{0}})_{∗} : π_{k}(F, v) → π_{k}(F ⟨S⟩, l_{s_{0}}(v)) is an isomorphism for 0 ≤ k ≤ n.

Proof. This follows by [4, Prop. 1.6] from the fact that p ∘ l_{s_{0}} = id_{F}. □

Lemma 4.2. Let F ⊆ O(V) satisfies the chain condition. Then |Link_{F}^{−}(v)| ≃ S^{|v|−2} for every v ∈ F.

Proof. Let v = (v_{1}, …, v_{n}). By definition Link_{F}^{−}(v) = {w ∈ F : w < v} = {(v_{i_{1}}, …, v_{i_{k}}) : k < n, i_{1} < ⋯ < i_{k}}. Hence |Link_{F}^{−}(v)| is isomorphic to the barycentric subdivision of the boundary of the standard simplex Δ_{n−1}. It is well known that ∂Δ_{n−1} ≃ S^{n−2}, hence |Link_{F}^{−}(v)| ≃ S^{|v|−2}. □

Theorem 4.3 (Nerve Theorem for Posets). Let V and T be two nonempty sets, F ⊆ O(V) and X ⊆ O(T). Assume X = ⋃_{v∈F} X_{v} such that if v ≤ w in F, then X_{w} ⊆ X_{v}. Let F, X and X_{v}, for every v ∈ F, satisfy the chain condition. Also assume

(i) for every v ∈ F, X_{v} is (l − |v| + 1)-acyclic (resp. (l − |v| + 1)-connected), (ii) for every x ∈ X, A_{x} := {v ∈ F : x ∈ X_{v}} is (l − |x| + 1)-acyclic (resp. (l − |x| + 1)-connected). Then H_{k}(F, ℤ) ≃ H_{k}(X, ℤ) for 0 ≤ k ≤ l (resp. F is l-connected if and only if X is l-connected).

Proof. Let F_{≤l+2} = {v ∈ F : |v| ≤ l + 2} and let i : F_{≤l+2} → F be the inclusion. Clearly |F_{≤l+2}| is the (l + 1)-skeleton of |F |, if we consider |F | as a cell complex whose k-cells are the |F_{≤v}| with |v| = k + 1. It is well known that i_{∗} : H_{k}(F_{≤l+2}, ℤ) → H_{k}(F, ℤ) and i_{∗} : π_{k}(F_{≤l+2}, v) → π_{k}(F, v) are isomorphisms for 0 ≤ k ≤ l (see [25], Chap. II, corollary 2.14, and [25], Chap. II, Corollary 3.10 and Chap. IV lemma 7.12.) So it is enough to prove the theorem for F_{≤l+2} and X_{≤l+2}. Thus assume F = F_{≤l+2} and X = X_{≤l+2}. We define Z ⊆ X × F as Z = {(x, v) : x ∈ X_{v}}. Consider the projections

f : Z → F, (x, v) ↦ v  ,  g : Z → X, (x, v) ↦ x.

First we prove that f^{−1}(v) ∼ v\f and g^{−1}(x) ∼ x\g, where ∼ means homotopy equivalence. By definition v\f = {(x, w) : w ≥ v, x ∈ X_{w}}. Define ϕ : v\f → f^{−1}(v), (x, w) ↦ (x, v). Consider the inclusion j : f^{−1}(v) → v\f. Clearly ϕ ∘ j(x, v) = ϕ(x, v) = (x, v) and j ∘ ϕ(x, w) = j(x, v) = (x, v) ≤ (x, w). So by 2.3(ii), v\f and f^{−1}(v) are homotopy equivalent. Similarly x\g ∼ g^{−1}(x). Now we prove that the maps f^{op} : Z^{op} → Y^{op} and g^{op} : Z^{op} → X^{op} satisfy the conditions of 3.6. First f^{op} : Z^{op} → Y^{op}; define the height function ht_{F^{op}} on F^{op} as ht_{F^{op}}(v) = l + 2 − |v|. It is easy to see that f^{op}/v ≃ v\f ∼ f^{−1}(v) ≃ X_{v}. Hence f^{op}/v is (l − |v| + 1)-acyclic (resp. (l − |v| + 1)-connected). But l−|v|+1 = (l+2−|v|)−1 = ht_{F^{op}}(v)−1, so f^{op}/v is (ht_{F^{op}}(v)−1)-acyclic (resp. (ht_{F^{op}}(v) − 1)-connected). Let n := l + 1. Clearly Link_{F^{op}}^{+}(v) = Link_{F}^{−}(v). By lemma 4.2, |Link_{F}^{−}(v)| is (|v|−3)-connected. But |v|−3 = l+1−(l+2−|v|)−2 = n − ht_{F^{op}}(v) − 2. Thus Link_{F^{op}}^{+}(v) is (n − ht_{F^{op}}(v) − 2)-acyclic (resp. (n − ht_{F^{op}}(v) − 2)-connected). Therefore by theorem 3.6, f_{∗} : H_{i}(Z, ℤ) → H_{i}(F, ℤ) is an isomorphism for 0 ≤ i ≤ l (resp. by 3.8, F is l-connected if and only if Z is l-connected). Now consider g^{op} : Z^{op} → X^{op}. We saw in the above that g^{op}/x ≃ x\g ∼ g^{−1}(x) and g^{−1}(x) = {(x, v) : x ∈ X_{v}} ≃ {v ∈ F : x ∈ X_{v}}. It is similar to the case of f^{op} to see that g^{op} satisfies the conditions of theorem 3.6, hence g_{∗} : H_{i}(Z, ℤ) → H_{i}(X, ℤ) is an isomorphism for 0 ≤ i ≤ l (resp. by 3.8, X is l-connected if and only if Z is l-connected). This completes the proof. □

Let K be a simplicial complex and {K_{i}}_{i∈I} a family of subcomplexes such that K = ⋃_{i∈I} K_{i}. The nerve of this family of subcomplexes of K is the simplicial complex N (K) on the vertex set I so that a finite subset σ ⊆ I is in N (K) if and only if ⋂_{i∈σ} K_{i} ≠ ∅. The nerve N (K) of K, with the inclusion relation, is a poset. As we already said we can consider a simplicial complex as a poset of its simplices.

Corollary 4.4 (Nerve Theorem). Let K be a simplicial complex and {K_{i}}_{i∈I} a family of subcomplexes such that K = ⋃_{i∈I} K_{i}. Suppose every nonempty finite intersection ⋂_{j=1}^{t} K_{i_{j}} is (l − t + 1)-acyclic (resp. (l − t + 1)connected). Then H_{k}(K, ℤ) ≃ H_{k}(N (K), ℤ) for 0 ≤ k ≤ l (resp. K is l-connected if and only if N (K) is l-connected).

Proof. Let V be the set of vertices of K. We give a total ordering to V and I. Put F = {(i_{1}, …, i_{r}) : i_{1} < ⋯ < i_{r} and ⋂_{j=1}^{r} K_{i_{j}} ≠ ∅} ⊆ O(I), X = {(x_{1}, …, x_{t}) : x_{1} < ⋯ < x_{t} and {x_{1}, …, x_{t}} is a simplex in K} ⊆ O(V) and for every (i_{1}, …, i_{r}) ∈ F, put X_{(i_{1},...,i_{r})} = {(x_{1}, …, x_{t}) ∈ X : {x_{1}, …, x_{t}} ∈ ⋂_{j=1}^{r} K_{i_{j}}}. It is not difficult to see that F ≃ N (K) and X ≃ K. Also one should notice that A_{x} := {v ∈ F : x ∈ X_{v}} is contractible for x ∈ X. We leave the details to interested readers. □

Remark 4.5. In [7], a special case of the theorem 4.3 is proved. The nerve theorem for a simplicial complex 4.4, in the stated generality, is proved for the first time in [3], see also [2, p. 1850]. For more information about different types of nerve theorem and more references about them see [2, p. 1850].

Lemma 4.6. Let F ⊆ O(V) satisfy the chain condition and let G : F^{op} → \underline{Ab} be a functor. Then the natural map ψ : ⨁_{v∈F, |v|=1} G(v) → H_{0}(F^{op}, G) is surjective.

Proof. By definition C_{0}(F^{op}, G) = ⨁_{v∈F^{op}} G(v), C_{1}(F^{op}, G) = ⨁_{v<v′∈F^{op}} G(v) and we have the chain complex

⋯ → C_{1}(F^{op}, G) →^{∂_{1}} C_{0}(F^{op}, G) → 0,

where ∂_{1} = d_{0}^{1} − d_{1}^{1}. Again by definition H_{0}(F^{op}, G) = C_{0}(F^{op}, G)/∂_{1}. Now let w ∈ F and |w| ≥ 2. Then there is a v ∈ F, v ≤ w, with |v| = 1. So w < v in F^{op}, and we have the component ∂_{1}|_{G(w)} : G(w) → G(w) ⊕ G(v), x ↦ d_{0}^{1}(x)−d_{1}^{1}(x) = d_{0}^{1}(x)−x. This shows that G(w) ⊆ im∂_{1} +imψ. Therefore H_{0}(F^{op}, G) is generated by the groups G(v) with |v| = 1. □

Theorem 4.7. Let V and T be two nonempty sets, F ⊆ O(V) and X ⊆ O(T). Assume X = ⋃_{v∈F} X_{v} such that if v ≤ w in F, then X_{w} ⊆ X_{v} and let F, X and X_{v}, for every v ∈ F, satisfy the chain condition. Also assume

(i) for every v ∈ F, X_{v} is min{l − 1, l − |v| + 1}-connected, (ii) for every x ∈ X, A_{x} := {v ∈ F : x ∈ X_{v}} is (l − |x| + 1)-connected, (iii) F is l-connected. Then X is (l − 1)-connected and the natural map

⨁_{v∈F, |v|=1}(i_{v})_{∗} : ⨁_{v∈F, |v|=1} H_{l}(X_{v}, ℤ) → H_{l}(X, ℤ)

is surjective, where i_{v} : X_{v} → X is the inclusion. Moreover, if for every v with |v| = 1, there is an l-connected Y_{v} with X_{v} ⊆ Y_{v} ⊆ X, then X is also l-connected.

Proof. If l = −1, then everything is easy. If l = 0, then for v of length one, X_{v} is nonempty, so X is nonempty. This shows that X is (−1)-connected. Also, every connected component of X intersects at least one X_{w} and therefore also contains a connected component of an X_{v} with |v| = 1. This gives the surjectivity of the homomorphism

⨁_{v∈F, |v|=1}(i_{v})_{∗} : ⨁_{v∈F, |v|=1} H_{0}(X_{v}, ℤ) → H_{0}(X, ℤ).

Now assume that, for every v of length one, X_{v} ⊆ Y_{v} where Y_{v} is connected. We prove, in a combinatorial way, that X is connected. Let x, y ∈ X, x ∈ X_{(v_{1})} and y ∈ X_{(v_{2})} where (v_{1}), (v_{2}) ∈ F. Since F is connected, there is a sequence (w_{1}), …, (w_{r}) ∈ F such that they give a path, in F, from (v_{1}) to (v_{2}), that is

(v_{1})  (w_{1})  (w_{r})  (v_{2})
⧹  ⧸  …  ⧹  ⧸
(v_{1}, w_{1})  (w_{r}, v_{2})  .

Since Y_{(v_{1})} is connected, x ∈ X_{(v_{1})} ⊆ Y_{(v_{1})} and X_{(v_{1},w_{1})} ≠ ∅, there is an element x_{1} ∈ X_{(v_{1},w_{1})} such that there is a path, in Y_{(v_{1})}, from x to x_{1}. Now x_{1} ∈ Y_{(w_{1})}. Similarly we can find x_{2} ∈ X_{(w_{1},w_{2})} such that there is a path, in Y_{(w_{1})}, from x_{1} to x_{2}. Now x_{2} ∈ Y_{(w_{2})}. Repeating this process finitely many times, we find a path from x to y. So X is connected. Hence we assume that l ≥ 1. As we said in the proof of theorem 4.3, we can assume that F = F_{≤l+2} and X = X_{≤l+2} and we define Z, f and g as we defined them there. Define the height function ht_{F^{op}} on F^{op} as ht_{F^{op}}(v) = l + 2 − |v|. As we proved in the proof of theorem 4.3, f^{op}/v ≃ v\f ∼ f^{−1}(v) ≃ X_{v}. Thus f^{op}/v is (ht_{F^{op}}(v) − 1)-connected if |v| > 1 and it is (ht_{F^{op}}(v) − 2)-connected if |v| = 1 and also |Link_{F^{op}}^{+}(v)| is (l + 1 − ht_{F^{op}}(v) − 2)-connected. By theorem 3.1, we have the first quadrant spectral sequence

E_{p,q}^{2} = H_{p}(F^{op}, v ↦ H_{q}(f^{op}/v, ℤ)) ⇒ H_{p+q}(Z^{op}, ℤ).

For 0 < q ≤ ht_{F^{op}}(v) − 2, H_{q}(f^{op}/v, ℤ) = 0. Define G_{q} : F^{op} → \underline{Ab}, G_{q}(v) = H_{q}(f^{op}/v, ℤ). Then G_{q}(v) = 0 for ht_{F^{op}}(v) ≥ q + 2, q > 0. By lemma 3.5, H_{p}(F^{op}, G_{q}) = 0 for p ≤ l + 1 − (q + 2). Therefore E_{p,q}^{2} = 0 for p + q ≤ l − 1, q > 0. If q = 0, arguing similarly to the proof of theorem 3.6, we get E_{p,0}^{2} = 0 if 0 < p ≤ l − 1 and E_{0,0}^{2} = ℤ. Also by the fact that F^{op} is l-connected we get the surjective homomorphism H_{l}(F^{op}, v ↦ H̃_{0}(f^{op}/v, ℤ)) ↠ E_{l,0}^{2}. Since l ≥ 1, H̃_{0}(f^{op}/v, ℤ) = 0 for all v ∈ F^{op} with ht_{F^{op}}(v) ≥ 1 and so H_{l}(F^{op}, v ↦ H̃_{0}(f^{op}/v, ℤ)) = 0 by lemma 3.5. Therefore E_{l,0}^{2} = 0. Let

G′_{q} : F^{op} → \underline{Ab}, G′_{q}(v) = \begin{cases}0 & if ht_{F^{op}}(v) < l + 1 \\ H_{q}(f^{op}/v, ℤ) & if ht_{F^{op}}(v) = l + 1 \end{cases}

and

G′′_{q} : F^{op} → \underline{Ab}, G′′_{q}(v) = \begin{cases}H_{q}(f^{op}/v, ℤ) & if ht_{F^{op}}(v) < l + 1 \\ 0 & if ht_{F^{op}}(v) = l + 1. \end{cases}

Then we have the short exact sequence 0 → G′_{q} → G_{q} → G′′_{q} → 0 and the associated long exact sequence

⋯ → H_{l−q}(F^{op}, G′_{q}) → H_{l−q}(F^{op}, G_{q}) →
H_{l−q}(F^{op}, G′′_{q}) → H_{l−q−1}(F^{op}, G′_{q}) → · · · .

If q > 0, then G′′_{q}(v) = 0 for 0 < q ≤ ht_{F^{op}}(v) − 1 and so by lemma 3.5, H_{p}(F^{op}, G′′_{q}) = 0 for p + q ≤ l, q > 0. Also if |v| = 1 then H_{0}(f^{op}/v, ℤ) = 0 for 0 < q ≤ ht_{F^{op}}(v) − 2 = l − 1. This shows G′_{q} = 0 for 0 < q ≤ l − 1. From the long exact sequence and the above calculation we get

E_{p,q}^{2} = \begin{cases}ℤ & if p = q = 0 \\ 0 & if 0 < p + q ≤ l, q ≠ l. \end{cases}

[figure] l + 1 ∗  E_{p,q}^{2}
l ∗ ∗
0 0 ∗
⋮ ⋮ ... ...
0 · · ... ∗
0 ℤ 0 ⋯ 0 0 ∗
0 1  l l + 1

Thus for 0 ≤ p+q ≤ l, q ≠ l, E_{p,q}^{2} ≃ ⋯ ≃ E_{p,q}^{∞} and there exist an integer r such that E_{0,l}^{2} ↠ ⋯ ↠ E_{0,l}^{r} ≃ E_{0,l}^{r+1} ≃ ⋯ ≃ E_{0,l}^{∞}. Hence we get a surjective map H_{0}(F^{op}, v ↦ H_{l}(f^{op}/v, ℤ)) ↠ H_{l}(Z^{op}, ℤ). By lemma 4.6, we have a surjective map ⨁_{v∈F, |v|=1} H_{l}(f^{op}/v, ℤ) ↠ H_{l}(Z^{op}, ℤ). Now consider the map g^{op} : Z^{op} → X^{op} and define the height function ht_{X^{op}}(x) = l + 2 − |x| on X^{op}. Arguing similarly to the proof of theorem 4.3 one sees that g_{∗} : H_{k}(Z, ℤ) → H_{k}(X, ℤ) is an isomorphism for 0 ≤ k ≤ l. Therefore we get a surjective map ⨁_{v∈F, |v|=1} H_{l}(X_{v}, ℤ) ↠ H_{l}(X, ℤ). We call it ψ. We prove that this map is the same map that we claimed. For v of length one consider the commutative diagram of posets

{v}^{op} ⟶ F^{op}
↗  ↗
f^{−1}({v})^{op} ⟶ Z^{op}
↘  ↘
X_{v}^{op} ⟶ X^{op}

By functoriality of the spectral sequence for the above diagram and lemma 4.6 we get the commutative diagram

H_{l}(f_{v}^{op}/v, ℤ)  ⟶^{(j_{v})_{∗}} ⨁_{v∈F,|v|=1} H_{l}(f^{op}/v, ℤ)
↓  ↓
H_{0}({v}^{op}, v ↦ H_{l}(f_{v}^{op}/v, ℤ)) ⟶ H_{0}(F^{op}, v ↦ H_{l}(f^{op}/v, ℤ))
↓  ↓
H_{l}(f^{−1}(v)^{op}, ℤ)  ⟶  H_{l}(Z^{op}, ℤ)
↓  ↓
H_{l}(X_{v}^{op}, ℤ)  ⟶^{(i_{v}^{op})_{∗}}  H_{l}(X^{op}, ℤ)

where j_{v} : f_{v}^{op}/v → f^{op}/v is the inclusion which is a homotopy equivalence as we already mentioned. It is not difficult to see that the composition of homomorphisms in the left column of the above diagram induces the identity map from H_{l}(X_{v}, ℤ), the composition of homomorphisms in the right column of above diagram induces the surjective map ψ and the last row induces the homomorphism (i_{v})_{∗}. This show that (i_{v})_{∗} = ψ|_{H_{l}(X_{v},ℤ)}. This completes the proof of surjectiveness. Now let for v of length one X_{v} ⊆ Y_{v} where Y_{v} is l-connected. Then we have the commutative diagram

\begin{matrix}H_{l}(X_{v}, ℤ) & ⟶^{(i_{v})_{∗}} & H_{l}(X, ℤ) \\ ↘ & ↗ \\ H_{l}(Y_{v}, ℤ) \end{matrix} .

By the assumption H_{l}(Y_{v}, ℤ) is trivial and this shows that (i_{v})_{∗} is the zero map. Hence by the surjectivity, H_{l}(X, ℤ) is trivial. If l ≥ 2, the nerve theorem 4.3 says that X is simply connected and by the Hurewicz theorem 2.1, X is l-connected. So the only case that is left is when l = 1. By theorem 3.8, X is 1-connected if and only if Z is 1-connected. So it is enough to prove that Z^{op} is 1-connected. Note that as we said, we can assume that F = F_{≤3} and X = X_{≤3}. Suppose F is a local system on Z^{op}. Define the functor G : F^{op} → \underline{Ab}, as

G(y) = \begin{cases}H_{0}(f^{op}/v, F) & if |v| = 1, 2 \\ H_{0}(Link_{F^{op}}^{+}(v), v′ ↦ H_{0}(f^{op}/v′, F)) & if |v| = 3 \end{cases} .

We prove that G is a local system on F^{op}. Put Z_{w} := g^{−1}(Y_{w}) for |w| = 1. If |v| = 1, 2, then f^{op}/v is 0-connected and f^{op}/v ⊆ Z_{w}^{op}, where w ≤ v, |w| = 1. By remark 3.9 we can assume that F = E ∘ g^{op} where E is a local system on X^{op}. Then F|_{Z_{w}^{op}} = E|_{Y_{w}^{op}} ∘ g^{op}|_{Z_{w}^{op}}. Since Y_{w}^{op} is 1-connected, E|_{Y_{w}^{op}} is a constant local system. This shows that F|_{Z_{w}^{op}} is a constant local system. So F|_{f^{op}/v} is a constant local system and since f^{op}/v is 0-connected we have H_{0}(f^{op}/v, ℤ) ≃ F(x), for every x ∈ f^{op}/v. If |v| = 3, with an argument similar to the proof of the theorem 3.8 and the above discussion one can get G(v) ≃ F(x) for every x ∈ f^{op}/v. This shows that G is a local system on F^{op}. Hence it is a constant local system, because F^{op} is 1-connected. It is easy to see that F ≃ G ∘ f. Therefore F is a constant system. Since X is connected by our homology calculation, by 3.7 we conclude that X is 1-connected. This completes the proof. □

5. Posets of unimodular sequences

Let R be an associative ring with unit. A vector (r_{1}, …, r_{n}) ∈ R^{n} is called unimodular if there exist s_{1}, …, s_{n} ∈ R such that Σ_{i=1}^{n}r_{i}s_{i} = 1, or equivalently if the submodule generated by this vector is a free summand of the left Rmodule R^{n}. We denote the standard basis of R^{n} by e_{1}, …, e_{n}. If n ≤ m, we assume that R^{n} is the submodule of R^{m} generated by e_{1}, …, e_{n} ∈ R^{m}. We say that a ring R satisfies the stable range condition (S_{m}), if m ≥ 1 is an integer so that for every unimodular vector (r_{0}, r_{1}, …, r_{m}) ∈ R^{m+1}, there exist t_{1}, …, t_{m} in R such that (r_{1} + r_{0}t_{1}, …, r_{m} + r_{0}t_{m}) ∈ R^{m} is unimodular. We say that R has stable rank m, we denote it with sr(R) = m, if m is the least number such that (S_{m}) holds. If such a number does not exist we say that sr(R) = ∞. Let U(R^{n}) denote the subposet of O(R^{n}) consisting of unimodular sequences. Recall that a sequence of vectors v_{1}, …, v_{k} in R^{n} is called unimodular when v_{1}, …, v_{k} is basis of a free direct summand of R^{n}. Note that if (v_{1}, …, v_{k}) ∈ O(R^{n}) and if n ≤ m, it is the same to say that (v_{1}, …, v_{k}) is unimodular as a sequence of vectors in R^{n} or as a sequence of vectors in R^{m}. We call an element (v_{1}, …, v_{k}) of U(R^{n}) a k-frame.

Theorem 5.1 (Van der Kallen). Let R be a ring with sr(R) < ∞ and n ≤ m+1. Let δ be 0 or 1. Then

(i) O(R^{n} + δe_{n+1}) ∩ U(R^{m}) is (n − sr(R) − 1)-connected. (ii) O(R^{n}+δe_{n+1})∩U(R^{m})_{v} is (n−sr(R)−|v|−1)-connected for all v ∈ U(R^{m}).

Proof. See [20, Thm. 2.6]. □

Example 5.2. Let R be a ring with sr(R) < ∞. Let n ≥ sr(R) + k + 1 and assume (v_{1}, …, v_{k}) ∈ U(R^{2n}). Set W = e_{2} + Σ_{i=2}^{n}Re_{2i}. Renumbering the basis one gets by theorem 5.1 that the poset F := O(W) ∩ U(R^{2n})_{(v_{1},...,v_{k})} is ((n−1)−sr(R)−k −1)-connected. Since n ≥ sr(R)+k +1, it follows that F is not empty. This shows that there is v ∈ W such that (v, v_{1}, …, v_{k}) ∈ U(R^{2n}). We will need such result in the next section but with a different method we can prove a sharper result. Compare this with lemma 5.4.

An n × k-matrix B with n < k is called unimodular if B has a right inverse. If B is an n × k-matrix and C ∈ GL_{k}(R), then B is unimodular if and only if CB

is unimodular. A matrix of the form ( 0 B^{1 u} ), where u is a row vector with

coordinates in R, is unimodular if and only if the matrix B is unimodular. We say that the ring R satisfies the stable range condition (S_{k}^{n}) if for every n × (n + k)-matrix B, there exists a vector r = (r_{1}, …, r_{n+k−1}) such that B ( 0 I_{n+k−1}^{1 r} ) = ( u B′ ), where the n × (n + k − 1)-matrix B′ is uni-

modular and u is the first column of the matrix B. Note that (S_{k}^{1}) is the same as (S_{k}).

Theorem 5.3 (Vaserstein). For every k ≥ 1 and n ≥ 1, a ring R satisfies (S_{k}) if and only if it satisfies (S_{k}^{n}).

Proof. The definition of (S_{k}^{n}) and the proof of this theorem is similar to the theorem [22, Thm. 3′] of Vaserstein. □

Lemma 5.4. Let R be ring with sr(R) < ∞ and let n ≥ sr(R)+k. Then for every (v_{1}, …, v_{k}) ∈ U(R^{2n}) there is a v ∈ e_{2} + Σ_{i=2}^{n}Re_{2i} such that (v, v_{1}, …, v_{k}) ∈ U(R^{2n}).

Proof. There is a permutation matrix A ∈ GL_{2n}(R) such that (e_{2} + Σ_{i=2}^{n}Re_{2i})A = e_{1} + Σ_{j=n+2}^{2n}Re_{j}. Let w_{i} = v_{i}A for i = 1, …, k. So (w_{1}, …, w_{k}) ∈ U(R^{2n}). Consider the k × 2n-matrix B whose i-th row is the vector w_{i}. By theorem 5.3 there exists a vector r = (r_{2}, …, r_{2n}) such that

B ( 0 I_{2n−1}^{1 r} ) = ( u_{1} B_{1} ), where the k × (2n − 1)-matrix B_{1} is unimod-

ular and u_{1} is the first column of the matrix B. Now let s = (s_{3}, …, s_{2n})

such that B_{1} ( 0 I_{2n_{2}}^{1 s} ) = ( u_{2} B_{2} ), where the k × (2n − 2)-matrix B_{2} is

unimodular and u_{2} is the first column of the matrix B_{1}. Now clearly

B ( 0 I_{2n−1}^{1 r} ) ( \begin{matrix}1 & 0 & 0 \\ 0 & 1 & s \\ 0 & 0 & I_{2n−2} \end{matrix} ) = ( u_{1} u_{2} B_{2} ) .

By continuing this process, n times, we find a 2n × 2n matrix C of the form

( \begin{matrix}1 & ∗ & ∗ & ∗ \\ & ^{.}^{.}. & ∗ & ∗ & ^{N} \\ & & 1 & ∗ & ∗ \\ & 0 & & 1 & ∗ \\ & & & & I_{n−1} \end{matrix} )

where N is an (n−1)×(n−1) matrix and BC = (L | M) where L is a k×(n+1) matrix and M is a unimodular k × (n − 1) matrix. Now let t = (t_{n+2}, …, t_{2n}) = −(first row of N). Then

\binom{1 0 … 0 t_{n+2} … t_{2n}}{B} C = \binom{1 ∗ ∗ 0 … 0}{∗ ∗ ∗ M} .

Since M is unimodular the right hand side of the above equality is unimodular. This shows that the matrix

\binom{1 0 … 0 t_{n+2} … t_{2n}}{B}

is unimodular. Put w = (1, 0, …, 0, t_{n+2}, …, t_{2n}). Then (w, w_{1}, …, w_{k}) ∈ U(R^{2n}). Now v = wA^{−1} is the one that we are looking for. □

6. Hyperbolic spaces and some posets

Let there be an involution on R, that is an automorphism of the additive group of R, R → R with r ↦ \overline{r}, such that \overline{\overline{r}} = r and \overline{rs} = \overline{s} \overline{r}. Let ϵ be an element in the center of R such that ϵ\overline{ϵ} = 1. Set R_{ϵ} := \{r − ϵ\overline{r} : r ∈ R\} and R^{ϵ} := \{r ∈ R : ϵ\overline{r} = −r\} and observe that R_{ϵ} ⊆ R^{ϵ}. A form parameter relative to the involution and ϵ is a subgroup Λ of (R, +) such that R_{ϵ} ⊆ Λ ⊆ R^{ϵ} and \overline{r}Λr ⊆ Λ, for all r ∈ R. Notice that R_{ϵ} and R^{ϵ} are form parameters. We denote them by Λ_{min} and Λ_{max}, respectively. If there is an s in the center of R such that s + \overline{s} ∈ R^{∗}, in particular if 2 ∈ R^{∗} , then Λ_{min} = Λ_{max}. Let e_{i,j}(r) be the 2n × 2n-matrix with r ∈ R in the (i, j) place and zero elsewhere. Consider Q_{n} = Σ_{i=1}^{n}e_{2i−1,2i}(1) ∈ M_{2n}(R) and F_{n} = Q_{n} + ϵ ^{t}Q_{n} = Σ_{i=1}^{n}(e_{2i−1,2i}(1) + e_{2i,2i−1}(ϵ)) ∈ GL_{2n}(R). Define the bilinear map h : R^{2n} × R^{2n} → R by h(x, y) = Σ_{i=1}^{n}(\overline{x_{2i−1}}y_{2i} + ϵ\overline{x_{2i}}y_{2i−1}) and q : R^{2n} → R/Λ by q(x) = Σ_{i=1}^{n}\overline{x_{2i−1}}x_{2i} mod Λ, where x = (x_{1}, …, x_{2n}), y = (y_{1}, …, y_{2n}) and \overline{x} = (\overline{x_{1}}, …, \overline{x_{2n}}). The triple (R^{2n}, h, q) is called a hyperbolic space. By definition the unitary group relative Λ is the group

U_{2n}^{ϵ}(R, Λ) := {A ∈ GL_{2n}(R) : h(xA, yA) = h(x, y), q(xA) = q(x), x, y ∈ R}.

For more general definitions and the properties of these spaces and groups see [8].

Example 6.1. (i) Let Λ = Λ_{max} = R. Then U_{2n}^{ϵ}(R, Λ) = {A ∈ GL_{2n}(R) : h(xA, yA) = h(x, y) for all x, y ∈ R^{2n}\} = \{A ∈ GL_{2n}(R) : ^{t}\overline{A}F_{n}A = F_{n}\}. In particular if ϵ = −1 and if the involution is the identity map id_{R}, then Λ_{max} = R. In This case U_{2n}^{ϵ}(R, Λ_{max}) := Sp_{2n}(R) is the usual symplectic group. Note that R is commutative in this case. (ii) Let Λ = Λ_{min} = 0. Then U_{2n}^{ϵ}(R, Λ) = {A ∈ GL_{2n}(R) : q(xA) = q(x) for all x ∈ R^{2n}}. In particular if ϵ = 1 and if the involution is the identity map id_{R}, then Λ_{min} = 0. In this case U_{2n}^{ϵ}(R, Λ_{min}) := O_{2n}(R) is the usual orthogonal group. As in the symplectic case, R is necessarily commutative. (iii) Let ϵ = −1 and the involution is not the identity map id_{R}. If Λ = Λ_{max} then U_{2n}^{ϵ}(R, Λ) := U_{2n}(R) is the classical unitary group corresponding to the involution.

Let σ be the permutation of the set of natural numbers given by σ(2i) = 2i − 1 and σ(2i − 1) = 2i. For 1 ≤ i, j ≤ 2n, i ≠ j, and every r ∈ R define

E_{i,j}(r) = \begin{cases}I_{2n} + e_{i,j}(r) & if i = 2k − 1, j = σ(i), r ∈ Λ \\ I_{2n} + e_{i,j}(r) & if i = 2k, j = σ(i), \overline{r} ∈ Λ \\ I_{2n} + e_{i,j}(r) + e_{σ(j),σ(i)}(−\overline{r}) & if i + j = 2k, i ≠ j \\ I_{2n} + e_{i,j}(r) + e_{σ(j),σ(i)}(−ϵ^{−1}\overline{r}) & if i ≠ σ(j), i = 2k − 1, j = 2l \\ I_{2n} + e_{i,j}(r) + e_{σ(j),σ(i)}(ϵ\overline{r}) & if i ≠ σ(j), i = 2k, j = 2l − 1 \end{cases}

where I_{2n} is the identity element of GL_{2n}(R). It is easy to see that E_{i,j}(r) ∈ U_{2n}^{ϵ}(R, Λ). Let EU _{2n}^{ϵ}(R, Λ) be the group generated by the E_{i,j}(r), r ∈ R. We call it elementary unitary group. A nonzero vector x ∈ R^{2n} is called isotropic if q(x) = 0. This shows automatically that if x is isotropic then h(x, x) = 0. We say that a subset S of R^{2n} is isotropic if for every x ∈ S, q(x) = 0 and for every x, y ∈ S, h(x, y) = 0. If h(x, y) = 0, then we say that x is perpendicular to y. We denote by ⟨S⟩ the submodule of R^{2n} generated by S, and by ⟨S⟩^{⊥} the submodule consisting of all the elements of R^{2n} which are perpendicular to all the elements of S. From now, we fix an involution, an ϵ, a form parameter Λ and we consider the triple (R^{2n}, h, q) as defined above.

Definition 6.2 (Transitivity condition). Let r ∈ R and define C_{r}^{ϵ}(R^{2n}, Λ) = {x ∈ Um(R^{2n}) : q(x) = r mod Λ}, where Um(R^{2n}) is the set of all unimodular vectors of R^{2n}. We say that R satisfies the transitivity condition (T_{n}), if EU _{2n}^{ϵ}(R, Λ) acts transitively on C_{r}^{ϵ}(R^{2n}, Λ), for every r ∈ R. It is easy to see that e_{1} + re_{2} ∈ C_{r}^{ϵ}(R^{2n}, Λ).

Definition 6.3 (Unitary stable range). We say that a ring R satisfies the unitary stable range condition (US_{m}) if R satisfies the conditions (S_{m}) and (T_{m+1}). We say that R has unitary stable rank m, we denote it with usr(R), if m is the least number such that (US_{m}) is satisfied. If such a number does not exist we say that usr(R) = ∞. Clearly sr(R) ≤ usr(R).

Remark 6.4. Our definition of unitary stable range is a little different than the one in [8]. In fact if (USR_{m+1}) satisfied then, by [8, Chap. VI, Thm. 4.7.1], (US_{m}) is satisfied where (USR_{m+1}) is the unitary stable range as defined in [8, Chap. VI, 4.6]. In comparison with the absolute stable rank asr(R) from [10], we have that if m ≥ asr(R) + 1 or if the involution is the identity map (so R is commutative) and m ≥ asr(R) then (US_{m}) is satisfied [10, 8.1].

Example 6.5. Let R be a commutative Noetherian ring where the dimension d of the maximal spectrum Mspec(R) is finite. If A is a finite R-algebra then usr(A) ≤ d + 1 (see [21, Thm. 2.8], [8, Thm. 6.1.4]). In particular if R is local ring or more generally a semilocal ring then usr(R) = 1 [8, 6.1.3].

Lemma 6.6. Let R be a ring with usr(R) < ∞. Assume n ≥ usr(R) + k and (v_{1}, …, v_{k}) ∈ U(R^{2n}). Then there is a hyperbolic basis {x_{1}, y_{1}, …, x_{n}, y_{n}} of R^{2n} such that v_{1}, …, v_{k} ∈ ⟨x_{1}, y_{1}, …, x_{k}, y_{k}⟩.

Proof. The proof is by induction on k. If k = 1, by definition of unitary stable range there is an E ∈ EU _{2n}^{ϵ}(R, Λ) such that v_{1}E = e_{1} + re_{2}. So the base of the induction is true. Let k ≥ 2 and assume the induction hypothesis. Arguing as in the base of the induction we can assume that v_{1} = (1, r, 0, …, 0), r ∈ R. Let W = e_{2} + Σ_{i=2}^{n}Re_{2i}. By lemma 5.4, choose w ∈ W so that (w, v_{1}, …, v_{k}) ∈ U(R^{2n}). Then (w, v_{1} − rw, v_{2}, …, v_{k}) ∈ U(R^{2n}). But (w, v_{1} − rw) is a hyperbolic pair, so there is an E ∈ EU _{2n}^{ϵ}(R, Λ) such that wE = e_{2n−1}, (v_{1} − rw)E = e_{2n} by [8, Chap. VI, Thm. 4.7.1]. Let (wE, (v_{1}−rw)E, v_{2}E, …, v_{k}E) =: (w_{0}, w_{1}, …, w_{k}) where w_{i} = (r_{i,1}, …, r_{i,2n}). Put u_{i} = w_{i} − r_{i,2n−1}e_{2n−1} − r_{i,2n}e_{2n} for 2 ≤ i ≤ k. Then (u_{2}, …, u_{k}) ∈ U(R^{2n−2}). Now by induction there is a hyperbolic basis {a_{2}, b_{2}, …, a_{n}, b_{n}} of R^{2n−2} such that u_{i} ∈ ⟨a_{2}, b_{2}, …, a_{k}, b_{k}⟩. Let a_{1} = e_{2n−1} and b_{1} = e_{2n}. Then w_{i} ∈ ⟨a_{1}, b_{1}, …, a_{k}, b_{k}⟩. But v_{1}E = w_{1} + rwE = e_{2n} + re_{2n−1}, v_{i}E = w_{i} for 2 ≤ i ≤ k and considering x_{i} = a_{i}E^{−1}, y_{i} = b_{i}E^{−1}, one sees that v_{1}, …, v_{k} ∈ ⟨x_{1}, y_{1}, …, x_{k}, y_{k}⟩. □

Definition 6.7. Let Z_{n} = {x ∈ R^{2n} : q(x) = 0}. We define the poset U′(R^{2n}) as U′(R^{2n}) := O(Z_{n}) ∩ U(R^{2n}).

Lemma 6.8. Let R be a ring with sr(R) < ∞ and n ≤ m. Then

(i) O(R^{2n}) ∩ U′(R^{2m}) is (n − sr(R) − 1)-connected, (ii) O(R^{2n})∩U′(R^{2m})_{v} is (n−sr(R)−|v|−1)-connected for every v ∈ U′(R^{2m}), (iii) O(R^{2n}) ∩ U′(R^{2m}) ∩ U(R^{2m})_{v} is (n − sr(R) − |v| − 1)-connected for every v ∈ U(R^{2m}).

Proof. Let W = ⟨e_{2}, e_{4}, …, e_{2n}⟩ and F := O(R^{2n}) ∩ U′(R^{2m}). It is easy to see that O(W)∩F = O(W)∩U(R^{2m}) and O(W)∩F_{u} = O(W)∩U(R^{2m})_{u} for every u ∈ U′(R^{2m}). By theorem 5.1, the poset O(W)∩F is (n−sr(R)−1)-connected and the poset O(W) ∩ F_{u} is (n − sr(R) − |u| − 1)-connected for every u ∈ F. It follows from lemma [20, 2.13 (i)] that F is (n − sr(R) − 1)-connected. The proof of (ii) and (iii) is similar to the proof of (i). □

Lemma 6.9. Let R be a ring with usr(R) < ∞ and let (v_{1}, …, v_{k}) ∈ U′(R^{2n}). If n ≥ usr(R)+k then O(⟨v_{1}, …, v_{k}⟩^{⊥})∩U′(R^{2n})_{(v_{1},...,v_{k})} is (n−usr(R)−k−1)- connected.

Proof. By lemma 6.6 there is a hyperbolic basis {x_{1}, y_{1}, …, x_{n}, y_{n}} of R^{2n} such that v_{1}, …, v_{k} ∈ ⟨x_{1}, y_{1}, …, x_{k}, y_{k}⟩. Let W = ⟨x_{k+1}, y_{k+1}, …, x_{n}, y_{n}⟩ ≃ R^{2(n−k)} and F := O(⟨v_{1}, …, v_{k}⟩^{⊥}) ∩ U′(R^{2n})_{(v_{1},...,v_{k})}. It is easy to see that O(W) ∩ F = O(W) ∩ U′(R^{2n}). Let V = ⟨v_{1}, …, v_{k}⟩, then ⟨x_{1}, y_{1}, …, x_{k}, y_{k}⟩ = V ⊕ P where P is a (finitely generated) projective module. Consider (u_{1}, …, u_{l}) ∈ F \O(W) and let u_{i} = x_{i} + y_{i} where x_{i} ∈ V and y_{i} ∈ P ⊕ W. One should notice that (u_{1} − x_{1}, …, u_{l} − x_{l}) ∈ U(R^{2n}) and not necessarily in U′(R^{2n}). It is not difficult to see that O(W) ∩ F_{(u_{1},...,u_{l})} = O(W) ∩ U′(R^{2n}) ∩ U(R^{2n})_{(u_{1}−x_{1},...,u_{l}−x_{l})}. By lemma 6.8, O(W) ∩ F is (n − k − usr(R) − 1)- connected and O(W) ∩ F_{u} is (n − k − usr(R) − |u| − 1)-connected for every u ∈ F \O(W). It follows from lemma [20, 2.13 (i)] that F is (n−usr(R)−k−1)- connected. □

7. Posets of isotropic and hyperbolic unimodular sequences

Let IU(R^{2n}) be the set of sequences (x_{1}, …, x_{k}), x_{i} ∈ R^{2n}, such that x_{1}, …, x_{k} form a basis for an isotropic direct summand of R^{2n}. Let HU(R^{2n}) be the set of sequences ((x_{1}, y_{1}), …, (x_{k}, y_{k})) such that (x_{1}, …, x_{k}), (y_{1}, …, y_{k}) ∈ IU(R^{2n}), h(x_{i}, y_{j}) = δ_{i,j}, where δ_{i,j} is the Kronecker delta. We call IU(R^{2n}) and HU(R^{2n}) the poset of isotropic unimodular sequences and the poset of hyperbolic unimodular sequences, respectively. For 1 ≤ k ≤ n, let IU(R^{2n}, k) and HU(R^{2n}, k) be the set of all elements of length k of IU(R^{2n}) and HU(R^{2n}) respectively. We call the elements of IU(R^{2n}, k) and HU(R^{2n}, k) the isotropic k-frames and the hyperbolic k-frames, respectively. Define the poset MU(R^{2n}) as the set of ((x_{1}, y_{1}), …, (x_{k}, y_{k})) ∈ O(R^{2n}×R^{2n}) such that, (i) (x_{1}, …, x_{k}) ∈ IU(R^{2n}), (ii) for each i, either y_{i} = 0 or (x_{j}, y_{i}) = δ_{ji}, (iii) ⟨y_{1}, …, y_{k}⟩ is isotropic. We identify IU(R^{2n}) with MU(R^{2n}) ∩ O(R^{2n} × {0}) and HU(R^{2n}) with MU(R^{2n}) ∩ O(R^{2n} × (R^{2n}\{0})).

Lemma 7.1. Let R be a ring with usr(R) < ∞. If n ≥ usr(R) + k then EU _{2n}^{ϵ}(R, Λ) acts transitively on IU(R^{2n}, k) and HU(R^{2n}, k).

Proof. The proof is by induction on k. If k = 1, by definition EU _{2n}^{ϵ}(R, Λ) acts transitively on IU(R^{2n}, 1) and by [8, Chap. VI, Thm. 4.7.1] the group EU _{2n}^{ϵ}(R, Λ) acts transitively on HU(R^{2n}, 1). The rest is an easy induction and the fact that for every isotropic k-frame (x_{1}, …, x_{k}) there is an isotropic k-frame (y_{1}, …, y_{k}) such that ((x_{1}, y_{1}), …, (x_{k}, y_{k})) is a hyperbolic k-frame [8, Chap. I, Cor. 3.7.4]. □

Lemma 7.2. Let R be a ring with usr(R) < ∞, and let n ≥ usr(R) + k. Let ((x_{1}, y_{1}), …, (x_{k}, y_{k})) ∈ HU(R^{2n}), (x_{1}, …, x_{k}) ∈ IU(R^{2n}) and V = ⟨x_{1}, …, x_{k}⟩. Then

(i) IU(R^{2n})_{(x_{1},...,x_{k})} ≃ IU(R^{2(n−k)})⟨V ⟩, (ii) HU(R^{2n}) ∩ MU(R^{2n})_{((x_{1},0),...,(x_{k},0))} ≃ HU(R^{2n})_{((x_{1},y_{1}),...,(x_{k},y_{k}))}⟨V × V ⟩, (iii) HU(R^{2n})_{((x_{1},y_{1}),...,(x_{k},y_{k}))} ≃ HU(R^{2(n−k)}).

Proof. See [5], the proof of lemma 3.4 and the proof of Thm. 3.2. □

For a real number l, by ⌊l⌋ we mean the largest integer n with n ≤ l.

Theorem 7.3. The poset IU(R^{2n}) is ⌊\frac{n−usr(R)−2}{2}⌋-connected and IU(R^{2n})_{x} is ⌊\frac{n−usr(R)−|x|−2}{2}⌋-connected for every x ∈ IU(R^{2n}).

Proof. If n ≤ usr(R), the result is clear, so let n > usr(R). Let X_{v} = IU(R^{2n})∩ U′(R^{2n})_{v} ∩ O(⟨v⟩^{⊥}), for every v ∈ U′(R^{2n}), and put X := ⋃_{v∈F} X_{v} where F = U′(R^{2n}). It follows from lemma 7.1 that IU(R^{2n})_{≤n−usr(R)} ⊆ X. So to treat IU(R^{2n}), it is enough to prove that X is ⌊\frac{n−usr(R)−2}{2}⌋-connected. First we prove that X_{v} is ⌊\frac{n−usr(R)−|v|−2}{2}⌋-connected for every v ∈ F. The proof is by descending induction on |v|. If |v| > n−usr(R), then ⌊\frac{n−usr(R)−|v|−2}{2}⌋ < −1. In this case there is nothing to prove. If n − usr(R) − 1 ≤ |v| ≤ n − usr(R), then ⌊\frac{n−usr(R)−|v|−2}{2}⌋ = −1, so we must prove that X_{v} is nonempty. This follows from lemma 6.6. Now assume |v| ≤ n − usr(R) − 2 and assume by induction that X_{w} is ⌊\frac{n−usr(R)−|w|−2}{2}⌋-connected for every w, with |w| > |v|. Let l = ⌊\frac{n−usr(R)−|v|−2}{2}⌋, and observe that n − |v| − usr(R) ≥ l + 2. Put T_{w} = IU(R^{2n})∩U′(R^{2n})_{wv} ∩O(⟨wv⟩^{⊥}) where w ∈ G_{v} = U′(R^{2n})_{v} ∩O(⟨v⟩^{⊥}) and put T := ⋃_{w∈G_{v}} T_{w} . It follows by lemma 6.6 that (X_{v})_{≤n−|v|−usr(R)} ⊆ T. So it is enough to prove that T is l-connected. The poset G_{v} is l-connected by lemma 6.9. By induction, T_{w} is ⌊\frac{n−usr(R)−|v|−|w|−2}{2}⌋-connected. But min\{l − 1, l − |w| + 1\} ≤ ⌊\frac{n−usr(R)−|v|−|w|−2}{2}⌋, so T_{w} is min\{l − 1, l − |w| + 1\}-connected. For every y ∈ T, A_{y} = {w ∈ G_{v} : y ∈ T_{w}} is isomorphic to U′(R^{2n})_{vy} ∩O(⟨vy⟩^{⊥}) so by lemma 6.9, it is (l − |y| + 1)-connected. Let w ∈ G_{v} with |w| = 1. For every z ∈ T_{w} we have wz ∈ X_{v}, so T_{w} is contained in a cone, call it C_{w}, inside X_{v}. Put C(T_{w}) = T_{w} ∪ (C_{w})_{≤n−|v|−usr(R)}. Thus C(T_{w}) ⊆ T. The poset C(T_{w}) is lconnected because C(T_{w})_{≤n−|v|−usr(R)} = (C_{w})_{≤n−|v|−usr(R)}. Now by theorems 5.1 and 4.7 , T is l-connected. In other words, we have now shown that X_{v} is ⌊\frac{n−usr(R)−|v|−2}{2}⌋-connected. By knowing this one can prove, in a similar way, that X is ⌊\frac{n−usr(R)−2}{2}⌋-connected. (Just pretend that |v| = 0.) Now consider the poset IU(R^{2n})_{x} for an x = (x_{1}, …, x_{k}) ∈ IU(R^{2n}). The proof is by induction on n. If n = 1, everything is easy. Similarly, we may assume n − usr(R) − |x| ≥ 0. Let l = ⌊\frac{n−usr(R)−|x|−2}{2}⌋. By lemma 7.2, IU(R^{2n})_{x} ≃ IU(R^{2(n−|x|)})⟨V ⟩, where V = ⟨x_{1}, …, x_{k}⟩. In the above we proved that IU(R^{2(n−|x|)}) is l-connected and by induction, the poset IU(R^{2(n−|x|)})_{y} is ⌊\frac{n−|x|−usr(R)−|y|−2}{2}⌋-connected for every y ∈ IU(R^{2(n−|x|)}). But l−|y| ≤ ⌊\frac{n−|x|−usr(R)−|y|−2}{2}⌋. So IU(R^{2(n−|x|)})⟨V ⟩ is l-connected by lemma 4.1. Therefore IU(R^{2n})_{x} is l-connected. □

Theorem 7.4. The poset HU(R^{2n}) is ⌊\frac{n−usr(R)−3}{2}⌋-connected and HU(R^{2n})_{x} is ⌊\frac{n−usr(R)−|x|−3}{2}⌋-connected for every x ∈ HU(R^{2n}).

Proof. The proof is by induction on n. If n = 1, then everything is trivial. Let F = IU(R^{2n}) and X_{v} = HU(R^{2n}) ∩ MU(R^{2n})_{v}, for every v ∈ F. Put X := ⋃_{v∈F} X_{v}. It follows from lemma 7.1 that HU(R^{2n})_{≤n−usr(R)} ⊆ X. Thus to treat HU(R^{2n}), it is enough to prove that X is ⌊\frac{n−usr(R)−3}{2}⌋- connected, and we may assume n ≥ usr(R) + 1. Take l = ⌊\frac{n−usr(R)−3}{2}⌋ and V = ⟨v_{1}, …, v_{k}⟩, where v = (v_{1}, …, v_{k}). By lemma 7.2, there is an isomorphism X_{v} ≃ HU(R^{2(n−|v|)})⟨V × V ⟩, if n ≥ usr(R) + |v|. By induction HU(R^{2(n−|v|)}) is ⌊\frac{n−|v|−usr(R)−3}{2}⌋-connected and again by induction HU(R^{2(n−|v|)})_{y} is ⌊\frac{n−|v|−usr(R)−|y|−3}{2}⌋-connected for every y ∈ HU(R^{2(n−|v|)}). So by lemma 4.1 , X_{v} is ⌊\frac{n−|v|−usr(R)−3}{2}⌋-connected. Thus the poset X_{v} is min{l − 1, l − |v| + 1}-connected. Let x = ((x_{1}, y_{1}), …, (x_{k}, y_{k})). It is easy to see that A_{x} = {v ∈ F : x ∈ X_{v}} ≃ IU(R^{2n})_{(x_{1},...,x_{k})}. By the above theorem 7.3, A_{x} is ⌊\frac{n−usr(R)−k−2}{2}⌋-connected. But l − |x| + 1 ≤ ⌊\frac{n−usr(R)−k−2}{2}⌋, so A_{x} is (l − |x| + 1)-connected. Let v = (v_{1}) ∈ F, |v| = 1, and let D_{v} := HU(R^{2n})_{(v_{1},w_{1})} ≃ HU(R^{2(n−1)}) where w_{1} ∈ R^{2n} is a hyperbolic dual of v_{1} ∈ R^{2n}. Then D_{v} ⊆ X_{v} and D_{v} is contained in a cone, call it C_{v}, inside HU(R^{2n}). Take C(D_{v}) := D_{v} ∪ (C_{v})_{≤n−usr(R)}. By induction D_{v} is ⌊\frac{n−1−usr(R)−3}{2}⌋-connected and so (l − 1)-connected. Let Y_{v} = X_{v} ∪ C(D_{v}). By the Mayer-Vietoris theorem and the fact that C(D_{v}) is l-connected, we get the exact sequence

H̃_{l}(D_{v}, ℤ) →^{(i_{v})_{∗}} H̃_{l}(X_{v}, ℤ) → H̃_{l}(Y_{v}, ℤ) → 0.

where i_{v} : D_{v} → X_{v} is the inclusion. By induction (D_{v})_{w} is ⌊\frac{n−1−usr(R)−|w|−3}{2}⌋- connected and so (l − |w|)-connected, for w ∈ D_{v}. By lemma 4.1(i) and lemma 7.2, (i_{v})_{∗} is an isomorphism, and by exactness of the above sequence we get H̃_{l}(Y_{v}, ℤ) = 0. If l ≥ 1 by the Van Kampen theorem π_{1}(Y_{v}, x) ≃ π_{1}(X_{v}, x)/N where x ∈ D_{v} and N is the normal subgroup generated by the image of the map (i_{v})_{∗} : π_{1}(D_{v}, x) → π_{1}(X_{v}, x). Now by lemma 4.1(ii), π_{1}(Y_{v}, x) is trivial. Thus by the Hurewicz theorem 2.1, Y_{v} is l-connected. By having all this we can apply theorem 4.7 and so X is l-connected. The fact that HU(R^{2n})_{x} is ⌊\frac{n−usr(R)−|x|−3}{2}⌋-connected follows from the above and lemma 7.2. □

Remark 7.5. One can define a more generalized version of hyperbolic space H(P) = P ⊕ P^{∗} where P is a finitely generated projective module. Charney in [5, 2.10] introduced the posets IU(P), HU(P) and conjectured that if P contains a free summand of rank on rank n then IU(P) and HU(P) are in fact highly connected. We leave it as exercise to the interested reader to prove this conjecture using the theorems 7.3 and 7.4 as in the proof of lemma 6.8. In fact one can prove that if P contains a free summand of rank n then IU(P) is ⌊\frac{n−usr(R)−2}{2}⌋-connected and HU(P) is ⌊\frac{n−usr(R)−3}{2}⌋-connected. Also, by assuming the high connectivity of the IU(R^{2n}), Charney proved that HU(R^{2n}) is highly connected. Our proof is different and relies on our theory, but we use ideas from her paper, such as the lemma 7.2 and her lemma 4.1, which is a modified version of work of Maazen [9].

8. Homology stability

From theorem 7.4 one can get the homology stability of unitary groups. The approach is well known.

Remark 8.1. To prove homology stability of this type one only needs high acyclicity of the corresponding poset, not high connectivity. But usually this type of posets are also highly connected. Here we also proved the high connectivity. In particular we wished to confirm the conjecture of Charney [5, 2.10], albeit with different bounds (see 7.5).

Theorem 8.2. Let R be a ring with usr(R) < ∞ and let the action of the unitary group on the Abelian group A is trivial. Then the homomorphism Inc_{∗} : H_{k}(U_{2n}^{ϵ}(R, Λ), A) → H_{k}(U_{2n+2}^{ϵ}(R, Λ), A) is surjective for n ≥ 2k + usr(R) + 2 and injective for n ≥ 2k + usr(R) + 3.

Proof. See [5, Section 4] and theorem 7.4. □

Remark 8.3. With the result of the previous section one also can prove homology stability of the unitary groups with twisted coefficients. For more information in this direction see [20, §5] and [5, 4.2].

References

[1] Betley, S. Homological stability for O_{n,n} over semi-local rings. Glasgow Math. J. 32 (1990), no. 2, 255–259.

[2] Björner, A. Topological methods. Handbook of combinatorics, Vol. 1, 2, 1819–1872, Elsevier, Amsterdam, (1995).

[3] Björner, A.; Lovász, L.; Vrećica, S. T.; Živaljević, R. T. Chessboard complexes and matching complexes. J. London Math. Soc. (2) 49 (1994), no. 1, 25–39.

[4] Charney R. On the problem of homology stability for congruence subgroups. Comm. in Algebra 12 (47) (1984), 2081–2123.

[5] Charney, R. A generalization of a theorem of Vogtmann. J. Pure Appl. Algebra 44 (1987), 107–125.

[6] Gabriel, P.; Zisman, M. Calculus of Fractions and Homotopy Theory. Band 35 Springer-Verlag New York, (1967).

[7] Grayson, D. R. Finite generation of K-groups of a curve over a finite field (after Daniel Quillen). Algebraic K-theory, Part I, 69–90, Lecture Notes in Math., 966, (1982).

[8] Knus, M. A. Quadratic and Hermitian forms over rings. Grundlehren der Mathematischen Wissenschaften , 294. Springer-Verlag, Berlin, 1991.

[9] Maazen, H. Homology stability for the general linear group. Thesis, Utrecht, (1979).

[10] Magurn, B.; Van der Kallen, W.; Vaserstein, L. Absolute stable rank and Witt cancellation for noncommutative rings, Invent. Math. 19 (1988), 525– 542.

[11] Milnor, J. The geometric realization of a semi-simplicial complex. Ann. of Math. (2) 65 (1957), 357–362.

[12] Panin, I. A. Homological stabilization for the orthogonal and symplectic groups. (Russian) Zap. Nauchn. Sem. Leningrad. Otdel. Mat. Inst. Steklov. (LOMI) 160 (1987), Anal. Teor. Chisel i Teor. Funktsii. 8, 222–228, 301–302 translation in J. Soviet Math. 52 (1990), no. 3, 3165–3170.

[13] Panin, I. A. On stabilization for orthogonal and symplectic algebraic Ktheory. (Russian) Algebra i Analiz 1 (1989), no. 3, 172–195 translation in Leningrad Math. J. 1 (1990), no. 3, 741–764.

[14] Quillen, D. Higher algebraic K-theory. I. Algebraic K-theory, I: Higher K-theories, pp. 85–147. Lecture Notes in Math., Vol. 341, Springer, Berlin 1973.

[15] Quillen, D. Finite generation of the groups K_{i} of rings of algebraic integers. Algebraic K-theory, I: Higher K-theories, pp. 179–198. Lecture Notes in Math., Vol. 341, Springer, Berlin, 1973.

[16] Quillen, D. Homotopy properties of the poset of nontrivial p-subgroups of a group. Adv. in Math. 28, no. 2 (1978), 101–128.

[17] Segal, G. Classifying spaces and spectral sequences. I.H.E.S. Publ. Math. No. 34 (1968), 105–112.

[18] Spanier, E. H. Algebraic Topology. McGram Hill (1966).

[19] Suslin, A. A. Stability in algebraic K-theory. Algebraic K-theory, Part I (Oberwolfach, 1980), pp. 304–333, Lecture Notes in Math., 966, Springer, Berlin-New York, 1982.

[20] Van der Kallen, W. Homology stability for linear groups. Invent. Math. 60 (1980), 269–295.

[21] Vaserstein, L. N. Stabilization of unitary and orthogonal groups over a ring with involution. (Russian) Mat. Sb. 81 (123) (1970) 328–351, translation in Math. USSR Sbornik 10 (1970), no. 3, 307–326.

[22] Vaserstein, L. N. The stable range of rings and the dimension of topological spaces. (Russian) Funkcional. Anal. i Prilozen. 5 (1971) no. 2, 17–27, translation in Functional Anal. Appl. 5 (1971), 102–110.

[23] Vogtmann, K. Spherical posets and homology stability for O_{n,n}. Topology 20 (1981), no. 2, 119–132.

[24] Weibel, C. A. An Introduction to Homological Algebra. Cambridge Studies in Advanced Mathematics, 38. Cambridge University Press, Cambridge, (1994).

[25] Whitehead, G. W. Elements of Homotopy Theory. Grad. Texts in Math. 61, Springer-Verlag, (1978).

Behrooz Mirzaii

Department of Mathematics Utrecht University P.O.Box 80.010 3508 TA Utrecht The Netherlands. mirzaii@math.uu.nl Wilberd van der Kallen Department of Mathematics Utrecht University P.O.Box 80.010 3508 TA Utrecht The Netherlands. vdkallen@math.uu.nl